Corner Market is considering adding a new product line that is expected to increase annual sales by $418,000 and cash expenses by $337,000. The initial investment will require $390,000 in fixed assets that will be depreciated using the 5-year MACRS method to a zero book value over the six-year life of the project. Ignore bonus depreciation. The company has a marginal tax rate of 21 percent. What is the value of the depreciation tax shield for the second year of this project ?

Group of answer choices

$13,650

$16,380

$26,208

$65,000

$81,900
Answer #1

Depreciation in Year 2 = 390,000*32% = $124,800 (using MACRS table)

Depreciation tax shield = Depreciation*Tax rate  

= 124,800*21%

= $26,208

Hence, the answer is $26,208